Sang invested some money at 8% interest. Sang also invested $184 more than 3 times that amount at 10%. How much is invested at each rate if Sang receives $1113.56 in interest after one year? (Round to two decimal places if necessary.)
Answer by graphmatics(170)

let x be the amount of money invested at 8% the 3*x+184 is the amount invested at 10%. The interest at 8% is .08*x. The interest at 10% is .10*(3*x+184). Since the sum of the interest is 1113.56 we get that
.08*x + .10*(3*x+184) = 1113.56
.08*x + .10*3*x + .10*184 = 1113.56
.38*x = 1113.56 - 18.4
x = 1131.96/.38
x = $2978.84
So the amount x invested at 8% is $2978.84 and the amount invested at 10% is
3*$2978.84 + $184 = $9120.56